An object of mass 4kg moves around a circle of radius 6m with a constant speed at 12m/s, calculate the angular acceleration


1
Expert's answer
2021-10-05T10:07:09-0400

"\varepsilon=\frac{v^2}{R^2}=\frac{12^2}{6^2}=4" s-2